anita is twice as old as anna. if 6 is subtracted from anna's age, and 2 is added to anita's age, anita will then be four times as old as anna. find their ages.
Answer by palanisamy(496)

Let the age of Anna be x.
Given,Anita is twice as old as Anna.
Then the age of Anitha = 2x
Also,if 6 is subtracted from anna's age and 2 is added to anita's age, anita will then be four times as old as anna
2x+2 = 4(x-6)
2x+2 = 4x-24
24+2 = 4x-2x
2x = 26
x = 13.
Therefore, the age of Anna is 13 and the age of Anitha is 26